OpenStudy (anonymous):

How do I do 2x^2=11x-5 Thanks!

OpenStudy (anonymous):

I think I got it... is it -10?

OpenStudy (ash2326):

We have \[2x^2=11x-5\] Do you know quadratic formula or do you know how to factor?

OpenStudy (anonymous):

yes I know both

OpenStudy (ash2326):

Nope, -10 is not the answer

OpenStudy (ash2326):

Try again

OpenStudy (anonymous):

Figured not. So I have to factor. I'll try tat.

OpenStudy (anonymous):

10

OpenStudy (anonymous):

So do I divide 2 on both sides and get x^2=5.5x-2.5 I think I messed up because I'm not sure how I would factor that.

OpenStudy (ash2326):

I'll explain you \[ 2x^2=11x-5\] Let's bring all the terms to one side \[2x^2-11x+5=0\] A standard quadratic is of the form \[ax^2+bx+c=0\] Tell me what's \( a\times b\) here?

OpenStudy (ash2326):

Sorry a*c?

OpenStudy (ash2326):

What's \(a \times c\) here?

OpenStudy (anonymous):

2 times 5 would be ten then

OpenStudy (anonymous):

I remember learning this, I hated it.

OpenStudy (ash2326):

Yeah:) What's b here?

OpenStudy (anonymous):

11, right

OpenStudy (ash2326):

It's very easy, After this you won't find trouble in solving these!!

OpenStudy (anonymous):

I hope so since my final is soon (:

OpenStudy (ash2326):

Uh-huh We need the sign also

OpenStudy (anonymous):

negative 11 then>

OpenStudy (ash2326):

Yeah :) Now I need you to split -11 into two numbers such that the product of the two numbers is 10 ( a*c)

OpenStudy (anonymous):

Hmm... I'm not sure. Sorry, 5.5 is my only guess but I know that's not right.

OpenStudy (anonymous):

I guess that would be -5.5

OpenStudy (ash2326):

It's easy let the two numbers be x and y, we have two conditions \[x+y=-11\] and \[x\times y=10\] First factor 10 and you'll get an idea

OpenStudy (anonymous):

I learned this as putting an equation that;s some thing like b=ax or some thing. But I'll try

OpenStudy (ash2326):

Yeah , just find the two no.s and it's almost done

OpenStudy (ash2326):

Will -1 and -10 work?

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Sorry, I was drawing a blank.

OpenStudy (anonymous):

So we stick those numbers some where... Into ax^2+bx+c ?

OpenStudy (ash2326):

\[2x^2-11x+5=0\] Now split -11x as -10x-1x \[2x^2 -10x-1x+5=0\]

OpenStudy (anonymous):

I never learned this way so sorry for my ignorance.

OpenStudy (ash2326):

Now factoring is easy \[2x^2 -10x-1x+5=0\] \[2x( x-5)-1(x-5)=0\] Now take (x-5) common, we get \[(x-5)(2x-1)=0\] Do you get this?

OpenStudy (ash2326):

It's fine:)

OpenStudy (anonymous):

So those are treated as two seperate things and factored semperately. SOrry for my typing, bad keyboard

OpenStudy (ash2326):

Yeah:) Can you find the values of x ?

OpenStudy (ash2326):

@randomemail98 I gotta go now, Sorry My friends will help you out @ParthKohli @Callisto Please help

OpenStudy (anonymous):

Alright bud

OpenStudy (anonymous):

Thank you so far

Parth (parthkohli):

Now, do you know the zero product rule?

OpenStudy (anonymous):

That a quadratic has to equal zero? If that's not it, then no.

Parth (parthkohli):

Okay, see :) Whenever (x - a)(x - b) [this is the factored quadratic expression] equals zero, then either x - a is 0, or x - b is 0.

Parth (parthkohli):

Now think of it like this. Whenever you multiply anything by 0, you get 0. Here, one of the numbers MUST be zero to make it 0.

OpenStudy (anonymous):

Yes, that makes sense

OpenStudy (anonymous):

The question is asking for 1 integer by the way

Parth (parthkohli):

Now, the factored form is \((x + 5)(2x + 1)\) Either x + 5 = 0 or 2x + 1 = 0 \( \color{Black}{\Rightarrow x + 5 = 0 }\) \( \color{Black}{\Rightarrow 2x + 1 = 0 }\) One of them doesn't have an integer as the solution.

OpenStudy (anonymous):

It says The integer is _____ Please don;t tell me I didn;t have to solve it or some thing, man. D:

OpenStudy (anonymous):

2x^2=11x-5 2x^2 -11x +5 =0 (2x-1)(x-5)=0

OpenStudy (anonymous):

its -11

OpenStudy (anonymous):

Callisto is smart lol

Parth (parthkohli):

Whenever you have \( \color{Black}{(x + a)(x + b) = 0 }\), then you make two different equations like this. \( \color{Black}{\Rightarrow x + a = 0 }\) \( \color{Black}{\Rightarrow x + b }\) This makes you determine the values.

Parth (parthkohli):

As it is \( \color{Black}{\Rightarrow (2x - 1)(x - 5) = 0 }\) Two equations: \( \color{Black}{\Rightarrow 2x - 1 = 0 }\) \( \color{Black}{\Rightarrow x - 5 = 0 }\) Solving, \( \color{Black}{\Rightarrow 2x = 1 }\) \( \color{Black}{\Rightarrow x = 0.5 }\) We're not done. \( \color{Black}{\Rightarrow x - 5 = 0 }\) \( \color{Black}{\Rightarrow x = 5 }\) :)

Parth (parthkohli):

The integer solution I see here is x = 5

OpenStudy (anonymous):

So oe of you say -11 and one say 5... D:

Parth (parthkohli):

Callisto, tell who's correct :D

OpenStudy (callisto):

@randomemail98 It's 5 ... '-11' means it's '-11x' in the question..

OpenStudy (anonymous):

Oh, I think I get it now. I;ve never seen it like that, but it's a good method. Thanks you guys, you didn;t have to all come here :P

OpenStudy (anonymous):

I'm gonna close this, thanks you all!